21) g(x) 2x-2 f(x)=x2 +3x Find (g f-2+x)
Answer #1

Given 961- 2x-2 fal= x²+2x Now find (gof) (-2+x) 96 f(-2+x)) =9[62+x)= 20 2-)] =9 [ <=4-44722-4] =9[x²= 2x = 2(x²2x) -2 = 2*2